At what net rate does heat radiate from a 244-m2 black roof on a night when the roof’s temperature is 20.34ºC and the surrounding temperature is 37.13 ºC? The emissivity of the roof is 0.379


1
Expert's answer
2021-03-24T19:44:09-0400

Convert temperature in degree Celsius to Kelvin:

T2=20.34+273.15=293.49 K,

T1=37.13+273.15=310.28 K.

The net rate of heat radiation is


"\frac{Q}{t}=A\epsilon\sigma(T_2^4-T_1^4),\\\space\\
\frac{Q}{t}=244·0.379·5.67·10^{-8}(293.49^4-310.28^4)=\\=-9700\text{ W}."